If you start contributing to your retirement account at the age of 18 with $25 a month and you increase your contributions by $5 a month per year, how much will you be contributing a month at the age of 40?

Answer by Theo(13342)

you start at age 18 with $25 a month contribution to the account.
each year you add $5 to the monthly contribution.
in year 1, the monthly contribution is 25.
in year 2, the monthly contribution is 30.
in year 3, the monthly contribution is 35.
etc.

this looks like an arithmetic progression type problem.

the formula for an arithmetic progression is An = A1 + (n-1)d

in this progression, A1 would be equal to 25
d would be equal to 5.
An would be 40 - 18 + 1 = 22.
when you are 18, it's year 1.
when you are 40, you add 40 - 18 = 22 to that to get year 23.

the formula becomes A23 = 25 + 22 * 5 = 135.

when you are 40, the monthly contribution will be 135 dollars.

this can be seen in the following excel printout.

$$$